Peggy had three times as many quarters as nickels. She had $1.60 in all. How many nickels and how many quarters did she have?
Which of the following equations could be used to solve the problem?

Answer by stanbon(75887)

Peggy had three times as many quarters as nickels. She had $1.60 in all. How many nickels and how many quarters did she have?
------------
Quantity Equation: q = 3n
Value Equation: 5n + 25q = 160 cents
------------------------------------------
Substitute for "q" and solve for "n":
5n + 25*3n = 160
5n + 75n = 160
80n = 160
n = 2 (# of nickels)
-------
Solve for "q":
q = 3n
q = 6 (# of quarters)
=====================
Cheers,
Stan H.
=====================